OpenStudy (anonymous):

How do you convert P=P(sub)0*a^t to P(sub)0*e^kt and vice versa? What's the difference?

OpenStudy (anonymous):

\[ P(t)=P_0\cdot a^t\to P_0e^{kt} \]

OpenStudy (anonymous):

Start with: \[ a^t=e^{kt} \]

OpenStudy (anonymous):

Solve for \(k\). This means you want to do logarithms. \[\begin{split} \ln(a^t)&=\ln(e^{kt})\\ t\ln(a)&=kt\\ \ln(a)&=k \end{split}\]

OpenStudy (anonymous):

Basically\[ P(t)=P_0\cdot a^t =P_0e^{\ln(a)t} \]

OpenStudy (anonymous):

Thanks so much!